Questionl The random variable X and Y have the following joint probability mass function: 0.14 0.27 0.2 0.1 0.03 0.15 0.1 a) Determine the b) Find P(X-Y>2). c) Find PX s3|Y20) d) Determine E(XY) e) Determine E(X) and E(Y). f) Are X and Y independent? marginal pmf for X and Y. Question 2 Let X and Y be independent random variables with pdf 2-y 0sxS 2 f(x)- f(p)- 0, otherwise 0, otherwise a) b) Find E(XY). Find Var (2X + 3Y).
Answer #1

1-a)

Following table shows the marginal pdfs of X and Y:

X
2 3 4 P(Y=y)
-1 0.1 0.2 0.05 0.35
Y 0 0.27 0.1 0.15 0.52
1 0 0.03 0.1 0.13
P(X=x) 0.37 0.33 0.3 1

b)

0.10.20.10.05 0.150.10.7

c)

P(Y > 0) P(У 0) + P(y-1) 0.52 + 0.13 0.65

P(Xleq 3cap Ygeq 0)=P(X=2,Y=0)+P(X=2,Y=1)+P(X=3,Y=0)+P(X=3,Y=1)=0.27+0+0.1+0.03=0.4

So the required probability is

=-= 200.650.6154

d)

Following table shows the calculations:

0.1 0.27 0 0.2 0.1 0.03 0.05 0.15 0.1 0.2 0.54 0 0 0 0 0.2 0.6 0.4 0.46 0 0 0 4 4 Total

e)

Following table shows the calculations:

P(X-x)Y 0.37 0.33 0.3 -1 0.35 0.52 0.13 0.74 0.99 1.2 2.93 0.35 0 0.13 0.22 0 4 Total

So we have

E(X)=sum xP(X=x)=2.93
y) =-0.22

f)

Since P(Y=-1) = 0.35, P(X=2) = 0.37 and P(X=2,Y=-1) = 0.1 so P(X=2, Y=-1) is not equal to P(X=2)*P(Y=-1). Therefore X and Y are not independent.